# Front Tracking for Scalar Conservation Laws with Spatially Heterogeneous Flux

### v2

@@ Abstract
In this article, we propose a novel front tracking scheme for scalar conservation laws with spatially heterogeneous, uniformly convex flux and demonstrate that the approximations converge to the entropy solution. The main tools we employ are Dafermos’ generalised characteristics and Kruzkov’s entropies.
@@ 1 Introduction
in $\mathbb{R}\times[0,\infty)$. Well-posedness has been studied for such equations when $f$ is “non-linear” in the conserved variable $u$ in some suitable sense [6]. Laws of this form are appropriate models, for example, for traffic flow with changing road conditions, such as varying maximum velocity. In this article, we focus on the case where $f$ is uniformly convex in $u$, and develop a front-tracking approach to the Cauchy problem (1). The same results hold for uniformly concave fluxes as well by a change of variables from $x$ to x-x.
A comprehensive treatment of conservation laws with spatially heterogeneous flux was conducted in [6], where the heterogeneity was assumed to be non-trivial in a compact domain. There, the flux was assumed to be ‘compactly non-homogeneous’, i.e. $f_{x}\equiv 0$ for $\mathinner{\!\left\lvert x\right\rvert}>K$ for some $K>0$. Convergence of a numerical scheme in this setting was explored previously using a finite volume scheme [24], discretising the heterogeneity and leveraging the theory of conservation laws with discontinuous flux [1]. Asymptotic emergence of simple shocks and their $L^{2}$ stability was analysed in [13].
Here, we deal with the heterogeneity directly in order to construct a sequence of approximate solutions, taking inspiration from Dafermos’ theory of generalised characteristics for scalar conservation laws with strictly convex flux [8] and the ‘front tracking’ method for scalar conservation laws more generally [7]. It was also implemented as a numerical method in [17]; see also the article by Holden and Holden [16].
@@
Standard reference texts for front tracking as applied to the Cauchy problem for systems of conservation laws in one spatial dimension include [4] and [18]. The results all generally assume a spatially homogeneous flux. Heterogeneity typically takes the form of sharp discontinuities in otherwise spatially homogeneous fluxes, and the front tracking method has been adapted to solve these Cauchy problem as well [14]. Such equations arise naturally in e.g. traffic flow, two-phase flows for oil extraction, etc. and a brief overview of the theory can be found in [18, Chapter 8]. Conservation laws with spatially discontinuous flux functions have been studied in several important papers; a necessarily incomplete list includes [1], [2], [9], [5], [11], [12], the references contained therein, and others.
@@
The structure of this article is as follows. Section 1.1 goes over some of the preliminary material for scalar conservation laws, including the entropy inequality for (1) and the theory of generalised characteristics. Section 1.2 outlines our structural assumptions on the flux. Section 2 details the front tracking algorithm for piecewise ‘stationary’ initial data and establishes a priori bounds, while Section 3 concludes well-posedness of the general Cauchy problem (1), i.e. existence of solutions satisfying entropy conditions (3), or equivalently (4). All functions of space and time are assumed to be càdlàg concerning the spatial variable, i.e. right-continuous with left limits, whenever the spatial traces exist unless otherwise specified.
@@ 1.1 Preliminaries
Classical solutions of (1) can be defined locally in time [19] for sufficiently regular initial data, but may break down in finite time. This motivates a notion of ‘weak solution’ that can be defined globally in time. The space-time divergence form of the equation naturally yields the following definition: $u\in L^{1}_{loc}([0,\infty);\mathbb{R})$ is a weak solution of the initial value problem (1) if, for all $\varphi\in C_{c}^{\infty}\left(\mathbb{R}\times[0,\infty)\right)$:
@@
$\begin{split}&\iint\mathinner{\!\left\lvert u(x,t)-k\right\rvert}\varphi_{t}+\operatorname{sgn}(u(x,t)-k)\left[f(x,u(x,t))-f(x,k)\right]\varphi_{x}dxdt\\ \geq&\iint\operatorname{sgn}(u(x,t)-k)f_{x}(x,k)\varphi(x,t)dxdt+\int\mathinner{\!\left\lvert u_{0}-k\right\rvert}\varphi(x,0)dx.\end{split}$ (4)
@@
[…]
@@ 2.1 Generalised Riemann problem
By construction, $u$ is a piecewise $C^{1}$ function that satisfies (1) classically away from the curve $(y(t),t)$ and satisfies the Rankine-Hugoniot condition along it. Moreover, at points of discontinuity, we have that $u(x-,t)=u_{l}(x)>u_{r}(x)=u(x+,t)$, hence $u$ is an entropy solution. Since $u$ also obeys a maximum principle by (UC), it is the unique entropy solution [20, Theorem 1].
@@ 3.1 A priori estimates
$\begin{split}&\iint\mathinner{\!\left\lvert u^{\delta}(x,t)-k\right\rvert}\varphi_{t}+\operatorname{sgn}(u^{\delta}(x,t)-k)\left[f^{\delta}(x,u^{\delta}(x,t))-f^{\delta}(x,k)\right]\varphi_{x}dxdt\\ \geq&\iint\operatorname{sgn}(u^{\delta}(x,t)-k)f_{x}^{\delta}(x,k)\varphi(x,t)dxdt+\int\mathinner{\!\left\lvert u_{0}-k\right\rvert}\varphi(x,0)dx.\end{split}$ (20)
@@ 3.2 Passing to the limit
$\geq$ $\iint\operatorname{sgn}(u(x,t)-k)f_{x}(x,k)\varphi(x,t)dxdt+\int\mathinner{\!\left\lvert u_{0}-k\right\rvert}\varphi(x,0)dx.$
Since $u\in L^{\infty}$ and $f$ satisfies (FSP), furthermore, we have uniqueness by [20, Theorem 1]. More precisely, we have $L^{1}$ stability with respect to a finite domain of dependence; bounded entropy solutions of (1) $u,v$ with respective initial values $u_{0},v_{0}$ satisfy

### v3

@@ Abstract
In this article, we propose a novel front tracking scheme for scalar conservation laws with spatially heterogeneous, uniformly convex flux and prove that approximations converge to the unique entropy solution. The main tools are Dafermos’ generalised characteristics and Kruzkov’s entropies. Crucially, our method handles fluxes where classical theory fails completely. As a concrete demonstration, we construct entropy solutions for a Cauchy problem with flux $f(x,u)=xu^{2}$, where bounded initial data can become unbounded in finite time, even on compact spatial domains. This finite-time blow-up violates the maximum principle, rendering all classical existence techniques—based on $L^{\infty}$ estimates and compactness—inapplicable. However, the flux $f(x,u(x,t))$ remains bounded despite $u$ blowing up, and our front tracking scheme exploits this to construct approximations that converge to an entropy solution.
@@ 1 Introduction
in $\mathbb{R}\times[0,\infty)$. An $L^{1}$ semigroup approach provided the initial breakthrough for such problems in the work of Kruzkov [25], who proved existence, uniqueness, and stability of ‘entropy’ solutions under some regularity assumptions on the flux. In this framework, however, although bounded solutions are unique when they exist, Kruzkov’s vanishing viscosity construction cannot guarantee the existence of solutions if the mixed partial derivative $f_{xu}$ is not uniformly bounded below. Even for the simplest convex heterogeneous fluxes such as $f(x,u)=(2+\sin x)u^{2}$, we have that $f_{xu}(x,u)=2u\cos x$, which has no uniform lower bound. The classical theory of scalar conservation laws can be found in [11].
Fluxes with spatial dependence have been analysed in the literature, such as on Riemannian or Lorentzian manifolds by LeFloch and Ben-Artzi in [4], motivated by geophysical and relativistic applications. Lengeler and Müller [26] proved the $L^{1}$ contraction property for scalar conservation laws on closed manifolds. The Cauchy problem was studied as a non-linear transport equation by Felix Otto in [27], who also derived regularising effect of the non-linearity. Colombo, Perrollaz, and Sylla [8] studied well-posedness for such equations when $f$ is “non-linear” in the conserved variable $u$, i.e. the set of zeroes of $f_{uu}$ is of measure zero, and the flux satisfies a super-linear growth assumption. This includes fluxes like $f(x,u)=(2+\sin x)u^{2}$. Convergence of a numerical scheme in this setting was explored by Abraham Sylla using a finite volume scheme [32], discretising the heterogeneity and leveraging the theory of conservation laws with discontinuous flux [1]. Asymptotic emergence of simple shocks and their $L^{2}$ stability was analysed by Ghoshal and Venkatesh in [17]. Recently, a mild generalisation of Kruzkov’s theory was obtained for scalar conservation laws in arbitrary dimension with spatial heterogeneity by Paz Hashah [20]. A purely $L^{1}$ approach, based on the Lions-Perthame-Tadmore kinetic formulation [28] was developed by Anne-Laure Dalibard for equations with heterogeneous flux in [12].
Laws of this form are also appropriate models, for example, for traffic flow with changing road conditions, such as varying maximum velocity. In this article, we focus on the case where $f$ is uniformly convex in $u$, and develop a front-tracking approach to the Cauchy problem (1). The same results hold for uniformly concave fluxes as well by a change of variables from $x$ to x-x.
We work with the heterogeneity directly in order to construct a sequence of approximate solutions, taking inspiration from Dafermos’ theory of generalised characteristics for scalar conservation laws with strictly convex flux [10] and the ‘front tracking’ method for scalar conservation laws more generally [9]. It was also implemented as a numerical method in [22]; see also the article by Holden and Holden [21].
@@
Standard reference texts for front tracking as applied to the Cauchy problem for systems of conservation laws in one spatial dimension include [6] and [23]. The results all generally assume a spatially homogeneous flux. Heterogeneity typically takes the form of sharp discontinuities in otherwise spatially homogeneous fluxes, and the front tracking method has been adapted to solve these Cauchy problem as well [18]. Such equations arise naturally in e.g. traffic flow, two-phase flows for oil extraction, etc. and a brief overview of the theory can be found in [23, Chapter 8]. Conservation laws with spatially discontinuous flux functions have been studied in several important papers; a necessarily incomplete list includes [1], [3], [13], [7], [15], [16], [2], the references contained therein, and others.
@@
The structure of this article is as follows. Section 1.1 goes over some of the preliminary material for scalar conservation laws, including the entropy inequality for (1) and the theory of generalised characteristics. Section 1.2 outlines our structural assumptions on the flux. Section 2 details the front tracking algorithm for piecewise ‘stationary’ initial data and establishes a priori bounds, while Section 3 concludes well-posedness of the general Cauchy problem (1), i.e. existence of solutions satisfying entropy conditions (3), or equivalently (4). In Section 4, we demonstrate the value of our method by proving the existence of entropy solutions with fluxes that are not contained in any of the existing theory.
All functions of space and time are assumed to be càdlàg concerning the spatial variable, i.e. right-continuous with left limits, whenever the spatial traces exist unless otherwise specified.
@@ 1.1 Preliminaries
The classical text [11] is a comprehensive reference for the literature on conservation laws. Here, we recall some of the basic concepts that we employ for our work. Differentiable solutions of (1) can be defined locally in time [24] for sufficiently regular initial data, but may break down in finite time. This motivates a notion of ‘weak solution’ that can be defined globally in time. The space-time divergence form of the equation naturally yields the following definition: $u\in L^{1}_{loc}([0,\infty);\mathbb{R})$ is a weak solution of the initial value problem (1) if, for all $\varphi\in C_{c}^{\infty}\left(\mathbb{R}\times[0,\infty)\right)$:
@@
$\begin{split}&\iint\mathinner{\!\left\lvert u(x,t)-k\right\rvert}\varphi_{t}+\operatorname{sgn}(u(x,t)-k)\left[f(x,u(x,t))-f(x,k)\right]\varphi_{x}dxdt\\ \geq&\iint\operatorname{sgn}(u(x,t)-k)f_{x}(x,k)\varphi(x,t)dxdt-\int\mathinner{\!\left\lvert u_{0}-k\right\rvert}\varphi(x,0)dx.\end{split}$ (4)
@@
on their respective domains, where $u(x\pm,t)$ respectively denote the left and right traces in space of $u$ at $(x,t)$. Since $f$ is convex in the second variable, $f_{u}$ is monotonically increasing, and entropy solutions of (1) satisfy the inequality $u_{-}\geq u_{+}$, hence the interval in (8) is well-defined for all $(x,t)$. Forward characteristics are also defined for points of the form $(x,0)$, but they may not be unique. A simple computation tells us that $f$ must be conserved along the trajectories of (6)
@@ 2.1 Generalised Riemann problem
By construction, $u$ is a piecewise $C^{1}$ function that satisfies (1) classically away from the curve $(y(t),t)$ and satisfies the Rankine-Hugoniot condition along it. Moreover, at points of discontinuity, we have that $u(x-,t)=u_{l}(x)>u_{r}(x)=u(x+,t)$, hence $u$ is an entropy solution. Since $u$ also obeys a maximum principle by (UC), it is the unique entropy solution [25, Theorem 1].
@@ 3.1 A priori estimates
$\begin{split}&\iint\mathinner{\!\left\lvert u^{\delta}(x,t)-k\right\rvert}\varphi_{t}+\operatorname{sgn}(u^{\delta}(x,t)-k)\left[f^{\delta}(x,u^{\delta}(x,t))-f^{\delta}(x,k)\right]\varphi_{x}dxdt\\ \geq&\iint\operatorname{sgn}(u^{\delta}(x,t)-k)f_{x}^{\delta}(x,k)\varphi(x,t)dxdt-\int\mathinner{\!\left\lvert u_{0}-k\right\rvert}\varphi(x,0)dx.\end{split}$ (20)
@@ 3.2 Passing to the limit
$\geq$ $\iint\operatorname{sgn}(u(x,t)-k)f_{x}(x,k)\varphi(x,t)dxdt-\int\mathinner{\!\left\lvert u_{0}-k\right\rvert}\varphi(x,0)dx.$
Since $u\in L^{\infty}$ and $f$ satisfies (FSP), furthermore, we have uniqueness by [25, Theorem 1]. More precisely, we have $L^{1}$ stability with respect to a finite domain of dependence; bounded entropy solutions of (1) $u,v$ with respective initial values $u_{0},v_{0}$ satisfy
@@
Finally, note that for fluxes which are merely strictly convex, we can add a term $\varepsilon u^{2}$. Then, we can solve the Cauchy problem by front tracking for $f_{\varepsilon}(x,u)=f(x,u)+\varepsilon u^{2}$ which satisfies all our assumptions instead of our original flux function. If we let $\varepsilon\to 0$, then, we will obtain a solution to the original Cauchy problem. Thus, (UC) can be relaxed to strict convexity.
4 Application of the method
This front tracking method for conservation laws with a spatially heterogeneous flux can be used to prove the existence of entropy solutions in the case of fluxes for which all classical methods fail, primarily due to their reliance on the maximum principle, e.g. the compensated compactness technique of [8]. The purely $L^{1}$ approach of [12] via the kinetic formulation, on the other hand, requires that $f_{x}$, or more generally the spatial divergence in several dimensions, be a uniformly bounded function of its arguments.
Now, if $f(x,u)=xu^{2}$, then it is trivial to see that $f$ does not satisfy any of the conditions necessary to conclude the existence of entropy solutions from the existing theory; note that $f$ is also not ‘weakly genuinely non-linear’ in the sense of [8], since $f(0,\cdot)$ is uniformly zero. As we will show below, entropy solutions of (1) with $f(x,u)=xu^{2}$ do not even satisfy the maximum principle, hence finite volume schemes such as [32] also cannot be used to conclude the existence of entropy solution.
The closest paper in the literature is [27] which precisely deals with a multiplicative flux. However, the existence of entropy solutions is merely assumed and not proved here, since the paper primarily deals with regularising effects and not existence of solutions.
4.1 Unbounded solutions
Although this flux also violates our assumption (UC) at $x=0$, we can adapt the front tracking method to deal with this issue. First, let us demonstrate that a priori $L^{\infty}$ estimates genuinely fail for this equation. Consider the unbounded function
$u(x,t)=\begin{cases}-1/\sqrt{x}&\text{ if }x>\gamma(t),\\ 0&\text{ otherwise},\end{cases}$
where $\gamma(t)$ is the Lipschitz curve
$\gamma(t)=\begin{cases}(t-2)^{2}/4&\text{ if }t\leq 2,\\ 0&\text{ if }t>2.\end{cases}$
At least for $t<2$, it is trivial to see that $u(x,t)$ is an entropy solution of (1) with bounded initial data. It solves the PDE pointwise on either side of the curve of discontinuity $\gamma(t)$, which in turn satisfies the Rankine-Hugoniot condition with the right sign. Indeed, the initial data $u_{0}(x)=u(x,0)$ is essentially a generalised Riemann data of the form (11), and thus $u(x,t)$ for $t<2$ is also a front tracking solution.
We claim that $u(x,t)$ is also an entropy solution for all $t\geq 2$. It is enough to show that the stationary solution $u(x,t+2)$ is an entropy solution of (1), and furthermore it is enough to prove the entropy inequality for compactly supported non-negative test functions in the open upper half plane, since the function is trivially $L^{1}_{loc}$ continuous.
Thus, let $v(x,t)=u(x,t)$, and note that $v$ is a classical solution of (1) for $x\neq 0$. For any compactly supported $\varphi\geq 0$ and $k\in\mathbb{R}$, note that for all $\varepsilon>0$,
$\iint_{\mathbb{R}^{2}_{+}}\mathinner{\!\left\lvert v-k\right\rvert}\varphi_{t}+\operatorname{sgn}(v-k)\left\{\left[f(x,u)-f(x,k)\right]\varphi_{x}-f_{x}(x,k)\varphi\right\}dxdt$
$=$ $\int_{0}^{\infty}\int_{0}^{\varepsilon}\mathinner{\!\left\lvert v-k\right\rvert}\varphi_{t}+\operatorname{sgn}(v-k)\left\{\left[v^{2}-k^{2}\right]x\varphi_{x}-k^{2}\varphi\right\}dxdt$
$+\int_{0}^{\infty}\int_{\varepsilon}^{\infty}\mathinner{\!\left\lvert v-k\right\rvert}\varphi_{t}+\operatorname{sgn}(v-k)\left\{\left[v^{2}-k^{2}\right]x\varphi_{x}-k^{2}\varphi\right\}dxdt$
Since all the terms are integrable, the first term vanishes as $\varepsilon\to 0$. Recall however that $v$ is a classical solution of (1) for $x>0$, and thus, applying integration by parts on the second integral yields
$\int_{0}^{\infty}\int_{\varepsilon}^{\infty}\mathinner{\!\left\lvert v-k\right\rvert}\varphi_{t}+\operatorname{sgn}(v-k)\left\{\left[v^{2}-k^{2}\right]x\varphi_{x}-k^{2}\varphi\right\}dxdt$
$=$ $\int_{0}^{\infty}\operatorname{sgn}\left(k+1/\sqrt{\varepsilon}\right)\varepsilon\left(\dfrac{1}{\varepsilon}-k^{2}\right)\varphi(\varepsilon,t)dt.$
Now, given any fixed $k\in\mathbb{R}$, this integral is clearly non-negative for all $\varepsilon$ small enough. Hence, we obtain the entropy inequality by letting $\varepsilon\to 0$. Since $\varphi\in C_{c}^{\infty}(\mathbb{R}^{2}_{+})$ and $k\in\mathbb{R}$ were arbitrary, this proves that $v$, and thus $u$, is indeed an entropy solution.
We remark that a similar construction can be carried out, mutatis mutandis, for solutions supported in $x<0$. Indeed, this method of proof works for any generalised Riemann problem whose $L^{\infty}$ norm blows up. This demonstration also naturally leads into our front tracking approximation method for solving the Cauchy problem for the flux $f(x,u)=xu^{2}$. If $g(x,u_{0}(x))\in BV(\mathbb{R})$, then we can construct $\delta$-approximate front tracking solutions by setting $g_{0}=0$ in a small enough neighbourhood of $x=0$, discretising as usual otherwise, and solving for front interactions, with the only change being: fronts that hit $x=0$ simply remain there. With this one minor tweak, we can conclude the existence of entropy solutions for the Cauchy problem even for fluxes such as $f(x,u)=xu^{2}$.